# InterHub: A Naturalistic Trajectory Dataset with Dense Interaction for Autonomous Driving

### v1

@@ Methods
Versatile benchmarks for key autonomous driving tasks
Autonomous driving tasks such as motion prediction and planning are generally validated in naturalistic driving datasets. However, due to the sparsity of challenging scenarios occurring in the natural collection, the performance of corresponding solutions in critical scenarios can be veiled by uninteresting cases, which occupy the majority, making the statistical measurement deceptively promising. In this subsection, we reveil the performance downgrade of state-of-the-art motion predictor and planners when transfer from natural collection to dense interaction subset of the driving scenario.
@@
  • UQnet  [45], our previous work, attained leading performance in the INTERPRET Challenge, ranking first on the leader-board of Single-Agent TrackINTERPRET: Interaction-dataset-based prediction challenge, single-agent track, organized by ICCV 2021 Competition. Available at: https://challenge.interaction-dataset.com/leader-board/. Last accessed on September 4-th 2024.. By Utilizing a deep learning model based on two-dimensional histograms and incorporating deep ensemble techniques, UQnet is capable of quantifying both epistemic uncertainty and aleatoric uncertainty, thereby significantly enhancing the generalizability of the missing rate compared to its previous state-of-the-art.
@@
  • ContextVAE  [47] is a context-aware model designed for predicting vehicle trajectories, which integrates environmental and agent dynamics through dual attention mechanisms to produce accurate, socially aware, and map-compliant predictions in real time. Since the lyft dataset did not have a unified race leaderboard to refer to, ContextVAE is selected as baseline model due to its leading performance on lyft level 5 in current publications[48, 49].
@@
To comprehensively assess the performance of these models, the following metrics are applied:
  • minADEk.    The minimum average euclidean distance between the predicted trajectory and the actual trajectory over all time steps is characterized as:
    $\text{minADE}_{k}=\min_{k\in\{1,\ldots,K\}}\frac{1}{T}\sum_{t=1}^{T}\sqrt{(\hat{x}_{t}-x_{t}^{k})^{2}+(\hat{y}_{t}-y_{t}^{k})^{2}}$ (1)
  • minFDEk.    The minimum euclidean distance between the predicted and actual positions at the final predicted time step is characterized as:
    $\text{minFDE}_{k}=\min_{k\in\{1,\ldots,K\}}\sqrt{(\hat{x}_{T}-x_{T}^{k})^{2}+(\hat{y}_{T}-y_{T}^{k})^{2}}$ (2)
  • Miss Rate (MR).    The proportion of predicted trajectories that deviate beyond a predefined threshold from the actual trajectory at the final predicted time step is characterized as:
    $\text{MR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\min_{k\in\{1,\ldots,K\}}\sqrt{(\hat{x}_{T}-x_{T}^{k})^{2}+(\hat{y}_{T}-y_{T}^{k})^{2}}>\delta\right)$ (3)
@@
In this section, we present the results of the planning and control tasks within challenging interaction scenarios. These tasks were evaluated in the context of OnSite Autonomous Driving Algorithm Challenge (www.onsite.com.cn) from 2023 to 2024. Organized by Tongji University and co-hosted by the China National Intelligent Connected Vehicle Innovation Center, the challenge attracts participation from more than 70 universities throughout the world. It utilized real-world natural driving data to simulate background traffic flow, covering realistic and demanding driving tasks, including car following, lane change, ramp merging, etc.
We analyzed the challenge results to investigate the relationship between the superior algorithms’ performance and the interaction intensity of the scenarios. The task performance score is measured from mainly safety, efficiency, and comfort with the evaluation tool of the OnSite challenge. We assessed the average performance score of the top 10 planners from more than 300 teams across over 700 driving interaction scenarios from the challenge lead-board. The interaction intensity of a given scenario is measured by the Maximum Average Interaction Intensity MaxIntensity.
$\text{MaxIntensity}=\max_{c\in C,t\in(T_{\text{start}},T_{\text{end}})}\frac{\text{MSAA}^{t}_{c}}{\text{Num}(c)}$ (4)
Here:
  • $C$ denotes the set of all groups of interacting agents within that scenario.
  • $c$ specifies a particular group of interacting vehicles.
  • $t$ is a time point within the interacting interval $(T_{\text{start}},T_{\text{end}})$.
@@
As shown in Figure 9, a consistent pattern is observed that as the interaction intensity increases, the average performance scores of the algorithms tend to decrease. This decline is particularly noticeable across interaction intensity intervals categorized as [0, 0.5) m/s², [0.5, 1) m/s², and [1, 1.5) m/s². Within these intervals, the algorithms exhibit a substantial decrease in their scores. However, for scenarios where the interaction intensity exceeds 1.5 m/s², the number of available scenarios is limited. Consequently, the algorithmic scores in these high-intensity scenarios show greater variability, which reduces the reliability of the relationship between interaction intensity and average score within this range.
The performance of autonomous driving algorithms in various scenarios can be assessed through their average scores. A higher average score indicates that autonomous vehicles controlled by the algorithm performed driving tasks more smoothly in the scenario, suggesting a lower difficulty. In contrast, a lower average score suggests a higher scenario difficulty. Our analysis underscores that scenarios characterized by intense interactions present more significant challenges for autonomous driving planning and control algorithms.
Discussion
Naturalistic driving datasets serve as foundational resources for empirical studies related to AVs. This study concentrates on a specific yet highly pertinent subset of the naturalistic driving, the driving interactions. In this section, we discuss the steps taken in our work, aiming to provide insights and directions for future research.
Ease of use across datasets. The precision, reproducibility, and universality of InterHub contribute to establishing a standardized benchmark for driving interaction research. InterHub supports the exploration of interaction scenarios by providing tools for data extraction and analysis, thus promoting usability in multiple datasets. This standardization facilitates researchers in comparing interaction events from various datasets, enhancing both the reproducibility and universality of interaction studies.
Interaction-related challenges in autonomous driving tasks. The rarity of risky driving scenarios poses a significant challenge in AV validation, where ensuring comprehensive safety is paramount. Mining these long-tailed, risky scenarios is thus a primary focus for testing AVs in interactive environments. Highly interactive scenarios have demonstrated particular difficulty for key AV tasks, including behavior prediction and motion planning. Our findings suggest that verification of AVs under such highly interactive conditions can help address the "tail" of challenging cases that are critical for achieving robust safety guarantees.
Underestimated potentials regarding dataset usage The INTERACTION dataset is widely used in research focused on interaction due to its dense collection of driving interaction events. While INTERACTION dataset captures significant interactions, its overall volume and diversity are constrained by its limited size. In contrast, larger datasets, such as WOMD, contain more, yet sparsely distributed, interactions that can offer valuable insights. The InterHub toolkit is specifically designed to facilitate access to these diverse data resources, enabling researchers to conduct more in-depth analyses and draw more comprehensive conclusions by leveraging both dense and sparse interaction data.
Recommendations for Future Driving Dataset Development. Data collected from an AV-centric perspective offers a rich foundation for studying interactions between autonomous and human-driven vehicles. Nonetheless, the absence of pure human-to-human interaction scenarios limits studies focused exclusively on human driving behaviors, which are crucial for AV navigation in mixed traffic settings. Moreover, inconsistencies in map formats present a substantial barrier to cross-dataset research. In our work, we preserved lane-level details to ensure commonality across maps from different datasets; however, additional details—such as region types and speed limits—must still be retrieved from the original datasets using the scenario indices provided in InterHub. Promoting a unified map description format, such as Lanelet [51], would significantly alleviate these inconsistencies and streamline research efforts.
From autonomous driving to future traffic system. Dense driving interaction datasets also provide new opportunities for advancing traffic management. While mature simulation techniques have been developed to model macro-level human driving behaviors for traffic management and infrastructure design, the effects of AVs remain less well understood. InterHub presents a unique resource that captures AV-specific interactions, enabling research on the integration of AVs within broader traffic systems and their subsequent impacts. Understanding these AV-centric interactions is critical as we move toward traffic ecosystems comprising a mix of human-driven and autonomous vehicles.
@@ Discussion
The codes for the unification of the raw driving records and the extraction and analysis of interaction events are documented and publicly available at https://github.com/zxc-tju/InterHub. We provide mainly three tools to help users navigate InterHub:

### v2

@@ Methods
Versatile benchmark for key autonomous driving tasks
Autonomous driving tasks, such as motion prediction and planning, are typically validated in naturalistic driving datasets. However, due to the sparsity of challenging scenarios occurring in the natural collection, the performance of corresponding solutions in critical scenarios can be obscured by the predominance of less interesting cases. This imbalance can lead to deceptively promising statistical measures. In this subsection, we reveal the performance degradation of state-of-the-art motion predictors and planners when transitioning from naturalistic driving datasets to the dense interaction subset of driving scenarios, highlighting the value of InterHub in providing a versatile benchmark for ensuring robust safety guarantees of AVs.
@@
  • UQnet  [45], our previous work, attained leading performance in the INTERPRET Challenge, ranking first on the leader-board of Single-Agent TrackINTERPRET: Interaction-dataset-based prediction challenge, single-agent track, organized by ICCV 2021 Competition. Available at: https://challenge.interaction-dataset.com/leader-board/. Last accessed on September 4-th 2024.. Using a deep learning model based on two-dimensional histograms and incorporating deep ensemble techniques, UQnet is capable of quantifying both epistemic uncertainty and aleatoric uncertainty, thereby significantly enhancing the generalizability of the missing rate compared to its previous state-of-the-art.
@@
  • ContextVAE  [47] is a context-aware model designed for predicting vehicle trajectories, which integrates environmental and agent dynamics through dual attention mechanisms to produce accurate, socially aware, and map-compliant predictions in real time. Since the lyft dataset did not have a unified race leaderboard to refer to, ContextVAE is selected as the baseline model due to its leading performance on lyft level 5 in current publications[48, 49].
@@
To comprehensively assess the performance of these models, the following metrics are applied. For each scenario, the model generates $K$ joint predictions $\hat{\mathbf{S}}^{(k)}=\{\hat{\mathbf{s}}_{i,t}^{(k)},i\in(1,\dots,N),t\in(1,\dots,T)\},k\in(1,\dots,K)$, where predictions are made for $T$ future time steps for $N$ agents. In a similar manner, the ground truth is represented as $\mathbf{S}=\{\mathbf{s}_{i,t}\}$.
  • minADE    The mean of the L2 norm between the real trajectory and the predicted trajectory for all agents over all time steps is as follows:
    $\text{minADE}=\min_{k}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\|s_{i,t}-\hat{s}_{i,t}^{(k)}\|_{2}$ (1)
  • minFDE    The L2 norm between the real trajectory and the predicted trajectory for all agents at the final predicted time step is as follows:
    $\text{minFDE}=\min_{k}\frac{1}{N}\sum_{i=1}^{N}\|s_{i,T}-\hat{s}_{i,T}^{(k)}\|_{2}$ (2)
  • Miss Rate (MR)    The proportion of predicted trajectories that deviate beyond a predefined threshold from the actual trajectory at the final predicted time step is as follows:
    $\text{MR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\min_{k}\|s_{i,T}-\hat{s}_{i,T}^{(k)}\|_{2}<\delta\right)$ (3)
    where $\mathbb{I}$ is an indicator function, when the minimum distance between the predicted value $\hat{s}_{i,T}^{(k)}$ and the ground truth value $s_{i,T}$ is smaller than a threshold $\delta$, the indicator function returns 1; otherwise, it returns 0.
@@
In this subsection, we present the results of the planning and control tasks within challenging interaction scenarios. These tasks were evaluated as part of the OnSite Autonomous Driving Algorithm Challenge (www.onsite.com.cn) held from 2023 to 2024. Organized by Tongji University and co-hosted by the China National Intelligent Connected Vehicle Innovation Center, the challenge attracted over 300 teams from more than 70 universities and institutions worldwide. With a specific focus on interaction scenarios such as head-on, crossing path, and merging path cases, selective scenarios were incorporated into the competition to challenge candidate planners.
We evaluated the average performance scores of the top 10 planners across over 700 interaction scenarios from the challenge leaderboard. Their task records were analyzed to investigate the relationship between algorithm performance and the interaction intensity of the scenarios. Performance scores were measured based on three primary factors: safety (time to collision, collision occurrence, and road boundary violations), efficiency (task completion and duration), and comfort (longitudinal and lateral acceleration, jerk, and yaw rate), using the evaluation tool provided by the OnSite challenge. Interaction intensity in each scenario was quantified by the average $\text{MSAA}^{t}$ over the involved agents during the interaction period.
@@
As shown in Figure 9, a consistent pattern emerges: as interaction intensity increases, the average performance scores of the algorithms tend to decrease, and the proportion of algorithms that fail to complete the task rises. This indicates that higher interaction intensity introduces escalating challenges for planning algorithms.
Mining long-tailed, high-risk scenarios is crucial for testing autonomous vehicles. Our findings suggest that verifying AV performance under such highly interactive conditions is essential for addressing the "tail" of challenging cases, which are critical for ensuring robust safety guarantees. To this end, InterHub is proposed not only as a dataset containing dense interaction cases but also as a hub for identifying and quantifying interaction events in a manner that is universally applicable across various disciplines.
@@
The codes for the unification of the raw driving records and the extraction and analysis of interaction events are documented and publicly available at https://github.com/zxc-tju/InterHub. We provide three tools to help users navigate InterHub: